### v1

# The Double-Copy Root of Hawking Thermality

Abstract
The Hawking radiation spectrum from a collapsing null shell can be derived via the double copy of a simpler gauge theory calculation. Analyzing the non-abelian Yang-Mills root of this process, we demonstrate that the radiation spectrum is thermal in the color charge eigenvalue $\lambda$, not energy. Considering the $SU(N_{c})$ gauge theory in the large $N_{c}$ limit, we find the differential spectrum $\mathrm{d}N/\mathrm{d}\lambda$ is a product of the gravitationally familiar Planck-like factor and the color phase space density, modeled here as the Wigner semicircle from random matrix theory. This reveals that apparent energy thermality in gravity is the direct dual of charge thermality in its underlying non-abelian gauge theory.
I Introduction
Classical solutions in General Relativity (GR), such as the Schwarzschild metric, can be viewed as an infinite sum of tree-level graviton diagrams sourced by their own self-interaction [1]. This perturbative picture suggests an immense complexity that is miraculously resummed in the exact solution. The classical double copy [2, 3, 4] reveals this is no miracle, but a direct consequence of gravity’s structure as the relativistic quantum double copy [5, 6, 7, 8] of a non-abelian Yang-Mills (YM) theory.
This principle is clearly manifest in metrics of the Kerr-Schild form, $g_{\mu\nu}=\eta_{\mu\nu}+\phi_{g}k_{\mu}k_{\nu}$. For a null, geodesic vector $k_{\mu}$, the Einstein tensor linearizes, $G_{\mu\nu}[g]=G^{(1)}_{\mu\nu}[h]$, meaning the full non-linear field equations reduce to a free wave equation. The infinite tower of graviton interactions collapses, and the solution behaves as if generated by a single, simple propagator. This simplicity is a direct consequence of the double copy, which guarantees that every operator in the Einstein-Hilbert action is completely specified by YM operator data [9, 10].
This correspondence is realized in the YM root. There, a specific gauge choice allows a potential with a fixed color orientation, $A^{a}_{\mu}=c^{a}\phi k_{\mu}$, to also behave as a simple propagator; it becomes an exact classical solution because its non-linear self-interactions vanish identically via the color algebra ($f^{a_{1}a_{2}a_{3}}c^{a_{2}}c^{a_{3}}=0$ if $c$ is aligned along a fixed color direction). The Kerr-Schild metric is therefore not merely analogous to a double copy; it is the literal double copy of the Yang-Mills propagator in the presence of an abelianized source. The geometric properties that linearize Einstein’s equations are the precise dual of the gauge choice that reduces Yang-Mills to its fundamental propagator.
While the classical solution appears abelian, the underlying root theory of general relativity must be non-abelian Yang-Mills. The duality relies on the kinematic numerators of YM amplitudes; indeed, their cubic and quartic gauge self-interactions are sufficient to generate all operators in GR through the double copy. Attempting to double-copy QED yields linearized gravity only. It has a consistent metric interpretation but none of the self-interaction that is the hallmark of Einstein-Hilbert. The celebrated generation of Newton from Coulomb is possible only because their non-relativistic potentials are described solely by mediator exchange in two-to-two scattering. Einstein requires Yang-Mills.
In a remarkable paper [11] of this past year, Aoude, O’Connell, and Sergola recovered the apparently thermal spectrum of Hawking radiation by emphasizing the on-shell nature of the original [12] calculation, setting the stage for a modern S-matrix perspective. The physical foundation of their approach lies in the principles of quantum field theory on a dynamic background. For a collapsing shell, the initial vacuum state ($|0_{\text{in}}\rangle$) is not an eigenstate of the final Hamiltonian and evolves under the S-matrix into a superposition of states containing real, outgoing particles — a formal description of particle creation from vacuum fluctuations.
Ref. [11] reminds us that we can compute the spectrum of this created radiation by considering the evolution of a single probe state. This probe is a computational tool; we calculate its scattering amplitude to characterize its dynamical interaction with the background responsible for particle creation. The calculation begins by computing a three-point tree-level amplitude for a probe scattering against the Vaidya background. By exponentiating this result via the Lippmann-Schwinger equation, the authors capture the exact result within the eikonal limit [^1] The eikonal limit describes the dynamics of particles whose trajectories are nearly light-like and forward-scattered. These are precisely the kinematics of the modes that just skim the forming event horizon to become the observed Hawking radiation., which re-sums the leading soft contributions to all orders.
The crucial result of this calculation is a logarithmic eikonal phase, $\chi(v_{0})\propto\log(-v_{0})$. The argument, $v_{0}$, represents the probe trajectory’s initial time offset $v_{0}<0$ relative to the moment of collapse. This phase is not specific to the probe but is a universal imprint left on any quantum mode by the extreme time-delay near the forming horizon. It is this mathematical structure that encodes the mixing of positive and negative frequency modes (a Bogoliubov transformation) that defines particle creation. The spectrum derived from the probe’s phase shift is therefore the spectrum of the particles spontaneously created from the vacuum. Critically, the Kerr-Schild nature of the Vaidya metric provides a unique opportunity to study this process not through the weak-strong duality of holography, but through the weak-weak duality of the double copy.
Note added in preperation: During the final stages of preparing this letter, two papers [14, 15] appeared with coordinated release. These papers carefully calculate in the abelian (Maxwell) limit of this root-Vaidya setup, confirming it yields a non-thermal, Bremsstrahlung-like energy spectrum. Our work complements their analysis by considering the non-abelian YM root. We find that the apparent thermality hidden in the abelian energy spectrum re-emerges in the natural charge of the full root theory: color.
II Derivation of the Spectrum
Here we summarize the derivation of the radiation spectrum from the eikonal S-matrix, referring readers to [11] for a comprehensive treatment in the double-copy case. The central object is the overlap amplitude between an initial single-particle wavepacket state $|\psi\rangle$ and a final on-shell momentum eigenstate $|p^{\prime}\rangle$. In the eikonal limit of high-energy, low-momentum-transfer ($q=p^{\prime}-p\to 0$) scattering, this amplitude can be expressed as an integral over the initial particle’s trajectory.
The key step involves replacing the quantum gluon in the 3-point scalar-gluon vertex with the classical background field $A_{\mu}^{\rm ext}(q)$. The S-matrix element at leading order becomes an integral over momentum transfer $q$:
$\langle p^{\prime}|\hat{S}-1|\psi\rangle\propto\int\mathrm{d}^{4}q\,\delta(2p^{\prime}\!\cdot\!q)\,e^{-iq\cdot b_{0}}\,i\mathcal{A}_{3}(p\to p^{\prime}).$ (1)
Here, the impact parameter $b_{0}$ parameterizes the initial trajectory, and the delta function enforces the on-shell condition for the probe in the eikonal limit. This constraint reduces the momentum-space integral to an integral along the classical, light-like worldline of the probe, $x^{\mu}(\sigma)=b_{0}^{\mu}+2\sigma p^{\prime\mu}$:
$\int\mathrm{d}^{4}q\,\delta(2p^{\prime}\!\cdot\!q)\,e^{-iq\cdot b_{0}}A_{\mu}^{\rm ext}(q)=\int\mathrm{d}\sigma\,A_{\mu}^{\rm ext}(x(\sigma)).$ (2)
The full S-matrix in this limit is found by solving the Lippmann-Schwinger equation, which re-sums soft radiation into an exponentiated eikonal phase, $S\approx e^{i\chi}$. For the collapsing shell background, $A_{\mu}^{\rm ext}\propto\theta(v)/r\,k_{\mu}$, the physical setup is straightforward. The integral for the phase is non-zero only where the source is active ($v>0$), which sets a physical, $v_{0}$-dependent lower bound on the integration region for an inbound trajectory parameterized by an early time $v_{0}<0$. The evaluation of this regulated $1/r$ integral results in the universal logarithmic form:
$\chi(v_{0})$ $\propto\int_{\sigma_{\rm min}}^{\infty}\mathrm{d}\sigma\,\frac{1}{r(\sigma)}\propto\int_{-v_{0}/E}^{\infty}\frac{\mathrm{d}r}{E}\frac{1}{r}$
$=\beta\times\log(-v_{0}/\mu).$ (3)
Here $E$ is the energy of the incoming probe, $\beta\propto g~\mathcal{C}$ is a constant of proportionality containing all coupling and charge interaction, while the IR scale $\mu$ absorbs all dependence on the arbitrary upper cutoff.
The final radiation spectrum is obtained from the Fourier transform of this time-dependent phase factor with respect to the radiated energy. As energy transfer is negligable in the eikonal limit we identify the radiated energy with for the energy of the incoming probe, $E$,
$\mathcal{A}(E)=\int_{-\infty}^{0}\mathrm{d}v_{0}\,e^{iEv_{0}}e^{i\beta\log(-v_{0}/\mu)}.$ (4)
The integral for $\mathcal{A}(E)$ is a standard result that evaluates to a Gamma function, $|\mathcal{A}(E)|^{2}\propto E^{-2}|\Gamma(1+i\beta)|^{2}$. Using the identity $|\Gamma(1+ix)|^{2}=\pi x/\sinh(\pi x)$, the differential particle number spectrum takes its final form:
$\frac{\mathrm{d}N}{\mathrm{d}E}\propto\frac{1}{E^{2}}\frac{\pi\beta}{\sinh(\pi\beta)}.$ (5)
This result forms the basis of our analysis. The physics is entirely encoded in the dependence of the phase coefficient $\beta$. For the abelian case, $\beta$ is a constant. For gravity, in sharp contrast, $\beta\propto E$, which yields the familiar Planck-like spectrum.
Gravity’s thermal form requires two ingredients: the logarithmic phase itself, which is a universal consequence of the long-range potential in four dimensions, and the linear energy dependence of its coefficient. This energy dependence has a spectacularly simple origin in the double-copy construction of the 3-point vertex: the YM color factor is replaced by a second copy of the kinematic numerator, effectively promoting the interaction’s momentum dependence from linear to quadratic.
Can the thermal nature of gravity and all the mysteries therein truly stem from something as elementary as a tree-level substitution in the three-point amplitude for minimal coupling? Yes, because this thermal structure is already present in the gauge theory as will become apparent when we analyze in full non-abelian generality.
III The Non-Abelian Root-Vaidya
We now consider the full non-abelian structure of the root-Vaidya background, a collapsing shell of charge with a fixed [^2] The assumption of a fixed color orientation $c^{a}$ is essential on two fronts. First, it ensures the background is an exact solution by causing the non-linear Yang-Mills term to vanish (abelianization). Second, it simplifies the probe’s eikonal phase to a c-number. If the background’s color varied along the probe’s trajectory, this phase would become a non-trivial path-ordered exponential. The S-matrix would become a matrix that rotates the probe’s color, meaning an initial color eigenstate would not evolve into a final one. In such a case, the entire concept of describing the interaction by a single phase dependent on a final-state charge eigenvalue $\lambda$ breaks down. color orientation $c^{a}$:
$A^{a}_{\mu}(x)=c^{a}\frac{Q_{0}\,\theta(v)}{r}k_{\mu}.$ (6)
As established, this is an exact solution to the classical YM equations. We consider a massless scalar probe in the adjoint of SU($N_{c}$) scattering off this background. The eikonal S-matrix is given by the path-ordered exponential $S=\mathcal{P}\exp(ig\int p^{\mu}A_{\mu}^{a}T^{a}\mathrm{d}\sigma)$. Because the background’s color vector $c^{a}$ is constant along the probe’s worldline, the color operator at every point is simply $c^{a}T^{a}$. This means the matrices at different points on the worldline commute, and the path-ordering becomes trivial.
The S-matrix thus reduces to a simple exponential, $S=\exp\left(g(c^{a}T^{a})\int p^{\mu}A_{\mu}^{\rm sp}\mathrm{d}\sigma\right)$, where $A_{\mu}^{\rm sp}$ is the spacetime part of the background. For a probe prepared in an eigenstate of the interaction operator $c^{a}T^{a}$, its color state is unchanged and simply acquires a phase. This phase is proportional to the corresponding eigenvalue $\lambda$, leading to a phase coefficient that is linear in the charge:
$\beta(\lambda)=C\lambda,$ (7)
where here the constant $C$ consolidates the coupling $g$, the source charge $Q_{0}$, and other color normalization. The amplitude for emitting a particle with a specific color eigenvalue $\lambda$ and energy $E$ then has a squared magnitude
$|\mathcal{A}(E,\lambda)|^{2}\propto\frac{1}{E^{2}}\frac{\pi C\lambda}{\sinh(\pi C\lambda)},$ (8)
where the $1/E^{2}$ prefactor is characteristic of Bremsstrahlung, but the $\sinh$ term introduces a thermal dependence on the color charge $\lambda$.
IV The Spectrum of Radiated Color
To compute an inclusive spectrum, we must integrate over all available color channels. In the large $N_{c}$ limit [^3] A crucial feature of the color-kinematics duality [5] is that the kinematic algebra is universal, holding for any number of external particles. The color algebra of any specific, finite-rank gauge group is not; for any finite $N_{c}$, certain color factors vanish for specific channels (e.g., the four-point singlet in SU(2)). For the duality to hold universally, the kinematic algebra cannot be restricted to any particular SU($N_{c}$). Instead, it is dual to a generic algebraic structure satisfying only antisymmetry and the Jacobi identity. The large $N_{c}$ limit of SU($N_{c}$) provides one ideal realization of such a structure. Our motivation for this limit should not be taken as any desire to access the planar sector or to simplify color topology. Rather, it is to treat the color space as a stand-in for this generic algebra, ensuring a dense spectrum of non-vanishing color channels., the canonical framework for this density of states is Random Matrix Theory (RMT) [18, 19], which predicts the Wigner semicircle law:
$\rho(\lambda)\propto\sqrt{R^{2}-\lambda^{2}},$ (9)
where $R$ defines the radius of the eigenvalue distribution. The final differential spectrum of radiated color charge is then the product of the dynamical emission probability and this density of available states, integrated over all radiated energies:
$\frac{\mathrm{d}N}{\mathrm{d}\lambda}\propto\underbrace{\frac{C\lambda}{\sinh(\pi C\lambda)}}_{\text{Dynamical Factor}}\times\underbrace{\sqrt{R^{2}-\lambda^{2}}}_{\text{Phase Space Factor}}.$ (10)
The physical nature of these two components constitutes the core result of this work. We now discuss its implications.
V Analysis and Interpretation
Equation (10) reveals a competition between dynamics and the available phase space in shaping the radiation.
  • The dynamical factor is thermal, arising from the universal resummation of soft radiation from a $1/r$ potential. It exponentially suppresses the emission of particles with large color charge ($\lambda\gg 1/C$).
  • The phase space factor is the Wigner semicircle, the spectral density of color charge. It is maximal at the origin and exhibits algebraic (square–root) suppression toward the boundary ($\lambda\to R$).
The observable spectrum’s shape depends on the dimensionless ratio $C/R$. For weak coupling ($C\ll R$), the spectrum traces the Wigner semicircle. For strong coupling ($C\gg R$), the thermal factor dominates, producing a Planck-like spectrum truncated by the phase space boundary at $\lambda=R$, as shown in Fig. 1. We emphasize that strong coupling in the eikonal context refers to the semi-classical regime of a large eikonal phase where $C\lambda\gg 1$, not a breakdown of perturbation theory; the soft series is already resummed (c.f.  WKB).
Figure 1: The spectrum of radiated color charge, $\mathrm{d}N/\mathrm{d}\lambda$, for fixed radius $R=1$ and varying effective coupling $C$. As $C$ increases, the spectrum transitions from the phase-space-dominated Wigner semicircle to a dynamically-dominated thermal distribution.
While we have adopted the Wigner semicircle as a well-motivated model [^4] The Wigner semicircle nicely reflects the universal eigenvalue density of random Hermitian matrices with SU($N_{c}$) symmetry and captures the expected large-$N_{c}$ distribution of color eigenvalues independent of microscopic details. for the kinematic phase space, we should emphasize that the emergence of thermality does not depend on this specific choice. Rather, the thermal behavior arises from the separate dynamical factor, which we derived from first principles. The thermal dynamics thus stands in opposition to the density of states, independent of its precise form. This opposition — where the strong exponential suppression from the dynamics overwhelms the gentler algebraic fall off of the phase space — is a key physical prediction of the non-abelian root.
VI Conclusion
The double copy provides a rigorous map from non-abelian gauge theory to gravity, uniting classical solutions with quantum radiation. We have shown that when applied to the Hawking effect for a collapsing shell, the thermality of the radiation is already present in the single-copy YM theory, not in its energy spectrum, but rather in its color spectrum. This democratizes the puzzle of gravitational thermalization, rooting Planck in the universal dynamics of soft radiation — accessible in the phase space of gauge theory’s intrinsic charge: color.
This perspective offers a rich context to engage with the challenges of black hole statistical mechanics. To be clear, we do not view the goal as simply to derive the statistical properties of the color degrees of freedom. Rather we find a clarifying opportunity in understanding how these properties are lifted to a kinematic algebra that governs the unitary quantum evolution of spacetime, from which its apparent thermality emerges. We expect the double copy to provide a robust functional arena to sharpen the provocative link between the semi-classical bathing of a probe in soft radiation and the one-loop instability of the quantum vacuum.
Acknowledgements.
We would like to especially thank Dionysius Anninos, Alexander C. Edison, Juna A. Kollmeier, Nicolas H. Pavao, and Sai Sasank Chava for insightful and supportive discussions, as well as comments on an earlier draft. We are particularly grateful for the warm support and encouragement of Donal O’Connell and Chris White who we contacted after ref. [14] appeared on the arXiv. This work was supported by the DOE under contract DE-SC0015910, and by Northwestern University via the Amplitudes and Insight Group, Department of Physics and Astronomy, and Weinberg College of Arts and Sciences.